# Information theory for data-driven model reduction in physics and biology

### v2

@@ Appendix F Variational IB
[…]
$=\mathbb{E}_{x}D_{\text{KL}}(p(h|x)\lVert\hat{p}(h)).$ (60)
If $p(h|x)$ and $\hat{p}(h)$ are chosen properly, the Kullback-Leibler can be expressed analytically which enables gradients to be effectively computed. As in Ref. [41], we take a Gaussian ansatz for $p(h|x)$ and let the marginal $\hat{p}(h)$ be a spherical unit-variance Gaussian. More concretely, encoded variables $H_{t}$ are sampled from $p(H_{t}|X_{t})$ by computing
$h_{t}=f_{W}(x_{t})+\sigma_{W}(x_{t})\eta,$ (61)
@@
To bound the entire loss from above, we must bound $I(Y,H)$ from below. We do this using the noise-contrastive estimate of the mutual information introduced in Ref. [81]. This recasts the problem as one of distinguishing samples from the distributions $p(y|h)$ and $p(y)$. Given a batch of $B$ pairs $(y,h)$ and one particular value $h_{i}$, one asks what the probability is that a sample $y_{j}$ is from $p(y|h_{i})$ (is a positive sample) and not $p(y)$ (is a negative sample). This probability is
$p(y_{j}=\text{pos}|h_{i})=\frac{\frac{p(y_{i}|h_{i})}{p(y_{i})}}{\sum_{k}^{B}\frac{p(y_{k}|h_{i})}{p(y_{k})}},$ (62)
@@
$\mathbb{E}[-\log p(y_{j}=\text{pos}|h_{i})],$ (63)
where the expectation is taken over indices $i$ and $j$, is closely related to the mutual information $I(Y,H)$. In the limit of infinite samples $B\rightarrow\infty$, this quantity is given by
[…]
If one had access to the probabilities $p(y_{i}=\text{pos}|h)$ appearing in Eq. 63, the mutual information could thus be easily determined. One attempts to estimate these probabilities by introducing a variational ansatz $f(y,h)$ to approximate the density ratio $\frac{p(y|h)}{p(y)}$. Typically this $f$ is represented by a neural network. One can then obtain a bound on the mutual information by minimizing
$\mathcal{L}_{B}=\mathbb{E}\left[-\log\frac{f(y_{i}|h_{i})}{\sum_{k}^{B}f(y_{k}|h_{i})}\right],$ (65)
@@
$I_{\text{NCE}}(Y,H)=\log B-\mathcal{L}_{B}\leq I_{\text{true}}(Y,H).$ (66)
@@
$\mathcal{L}_{\text{VIB}}=D_{KL}(p(H_{t}|X_{t})\lVert\hat{p}(H_{t}))-\beta I_{\text{NCE}}(X_{t+\Delta t};H_{t}),$ (67)
@@
The VIB loss Eq. 67 is very similar to the loss function for a time-lagged $\beta$-variational autoencoder ($\beta$-VAE) [83], which have a form $\mathcal{L}=D_{KL}-\mathcal{L}_{\text{rec}}$, where the reconstruction term $\mathcal{L}_{\text{rec}}$ measures the deviation from the true future state $X_{t+\Delta t}$ and the reconstruction from the latent variable $H_{t}$. In the VIB, this term is replaced by the mutual information $I_{\text{NCE}}$: rather than searching for a latent variable which can reconstruct the While a standard $\beta$-VAE with mean squared error (MSE) loss finds a latent variable that can reconstruct the full state at a time $\Delta t$ in the future, the VIB merely tries to reconstruct the statistics of $X_{t+\Delta t}$ conditioned on $X_{t}$.
@@
Figure S6: Variational IB discovers slow variables. (a) Variational IB applied to the Lorenz system. Color corresponds the encoding value $h_{t}$. (b) Time correlation functions of the encoded variable $C_{XY}(\tau)=\langle(X(\tau)-\bar{X})(Y(0)-\bar{Y})\rangle$ for high compression (black) and no compression (red). Blue shows the decay of the subleading Koopman eigenfunction using numerically-obtained eigenvalues. (c-d) Lorenz attractor projected onto the $x$-$y$ plane, colored by encoding variable. (e) Same projection, colored by the value of the true subleading Koopman eigenfunction. (f) Dynamics of $h_{t}$ obtained by encoding the state at every time during the true trajectory. Top corresponds to no compression ($\beta=\infty$), bottom corresponds to high compression ($\beta=4000$).
@@
$\dot{\theta}$ $=\omega.$ (68)
Equation (68) is the normal form (in polar coordinates) for dynamics near a Hopf bifurcation [89]. For $\mu>0$ this system exhibits a circular limit cycle of radius $\sqrt{\mu}$ (Fig. S5b). In Section G we show that this system has infinite purely imaginary eigenvalues $\lambda_{n}=in\omega$ and that as a result, encoders which exactly encode the angle coordinate, $p(h|r,\theta)\propto\delta(h-\theta)$, are solutions of the IB optimization problem.
The exact IB calculation breaks down for perfectly deterministic dynamics, hence we add a small amount of white noise to the dynamics. While this slightly perturbs the spectral content of the transfer operator, we still find our results to be consistent with the those expected for the deterministic case (see Section G). As we increase the size of the encoding alphabet $N_{H}$, the encoder partitions space by finer and finer angular wedges. The learned encoding is invariant under permutations of the encoded symbols. Upon reordering, we find that, for large alphabets $N_{H}\gg 1$, the encoder indeed approximates $p(h|r,\theta)\propto\delta(h-\theta)$ as expected (Fig. S5c).
Using VIB, we learn a continuous $h_{t}$ which can be computed directly from samples of the state variable. The encoding $h_{t}$ learned by VIB closely approximates the angle coordinate (Fig. S5e-f) and is nearly uncorrelated with $r$ (inset). This shows that our mathematical results illustrated for exact IB in Sec. D hold also in the approximate framework of VIB.
@@
[…]
@@
We illustrate this with a prototypical example of deterministic chaotic dynamics, the Lorenz system [90]. In the steady state, the state variable of the Lorenz system resides on a chaotic attractor consisting of two “lobes” which encircle unstable fixed points (Fig. S6a). The encoding variable $h$ learned by VIB is the slowest varying function of the state $\mathbf{x}$, decaying as $\exp(\lambda^{\text{Re}}_{1}t)$ where $\lambda^{\text{Re}}_{1}$ is the real part of the first subleading eigenvalue of the transfer operator (here Perron-Frobenius operator) (Fig. S6b). Eigenvalues $\lambda_{n}$ as well as right eigenvectors $\phi_{n}(x)$ of the Perron-Frobenius operator were computed numerically using the Ulam method. The correspondence between $\phi_{1}(x)$ and $h$ disappears as $\beta$ is increased (Fig. S6c-e). The dynamics of $h$ also become notably less “slow”, and are instead nearly constant with occasional large jumps (Fig. S6f). This shows that in order for $h$ to be a valid slow variable, the compression term in Eq. (67) is crucial. These results are consistent with those obtained for exact IB, where we showed that the encoder incorporates the slow modes at low beta (high compression).
Appendix G Optimal encoding for Hopf normal form dynamics
$\dot{x}$ $=(\mu-x^{2}-y^{2})x-\omega y$ (69)
$\dot{y}$ $=(\mu-x^{2}-y^{2})y-\omega x$ (70)
Expressed in polar coordinates, we have
$f_{x}\hat{e}_{x}$ $=f_{x}\cos\theta\hat{e}_{r}-f_{x}\sin\theta\hat{e}_{\theta}$ (71)
$f_{y}\hat{e}_{y}$ $=f_{y}\sin\theta\hat{e}_{r}+f_{y}\cos\theta\hat{e}_{\theta}$ (72)
$f_{i}\partial_{i}\phi$ $=(\partial_{r}\phi)(\mu-r^{2})r+\frac{1}{r}(\partial_{\theta}\phi)\omega r$ (73)
$=r(\mu-r^{2})(\partial_{r}\phi)+\omega\partial_{\theta}\phi$ (74)
For the angle coordinate alone the situation is much simpler,
$\omega\partial_{\theta}\phi(\theta)=\lambda\phi(\theta)$
$\phi_{n}(\theta)=e^{in\theta},\quad\quad\lambda_{n}=in\omega.$
[…]
We next ask whether an encoder of the form $p(h|\theta)\propto\delta(h-\theta)$ is a solution. Recall that the probability distribution appearing in Eqn. 77 is a distribution over future positions, $p(R_{t+\Delta t}=r^{\prime},\Theta_{t+\Delta t}=\theta+\omega\Delta t|h)$. This distribution can be calculated as
[…]
$\propto\delta((\theta_{t+\Delta t}-\omega\Delta t)-h)$ (79)
where we used that the two terms in the integrand of the first equation are both delta functions. Plugging this into (77) shows that $p(h|\theta)\propto\delta(h-\theta)$ is consistent, and hence is an optimal encoding. In the presence of noise, the eigenvectors are perturbed and may gain a dependence on $r$ (see Fig. S7).
Appendix H VIB for fluid flow around a cylinder
@@
Here we provide some additional details on the computation used to generate components of Fig. 5 of the main text. The true Koopman modes were computed using dynamical mode decomposition (DMD) [11, 12, 91], also described below in Section K. DMD attempts to find a finite dimensional approximation of the Koopman operator using $n$ snapshots of the system’s state $x\in\mathbb{R}^{d}$ which are assembled into a data matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$. One then attempts to find a linear evolution operator $K$ which propagates the state forward in time
@@
$\phi_{n}[\mathbf{v}]=\langle\mathbf{v}(\mathbf{x}),\mathbf{m}^{(n)}(\mathbf{x})\rangle,$ (80)
where $\mathbf{m}^{(n)}$ is the $n$-th “Koopman mode” and angled brackets denote integration over space. To compute these modes and the corresponding transfer operator spectrum, we use dynamic mode decomposition (DMD; see SI Section H) [11, 12].
We allow VIB to learn a two-dimensional encoding variable $[h_{0},h_{1}]$, so that it can learn the complete first eigenfunction rather than only the real or imaginary part. For the purposes of comparing our learned variable with $\phi_{1}$, we construct a complex $h=h_{0}+ih_{1}$ out of the two learned components.
To understand which function has been learned by the neural network, we check whether the learned functions $h_{i}[\mathbf{v}]$ are of the form Eq. (80). This can be done by examining gradients of the network with respect to the input field
$\frac{\partial h}{\partial v_{j}}=m^{(\text{IB})}_{j}+g_{\text{res},j}(\mathbf{v}(\mathbf{x}))$ (81)
@@
$\frac{\partial\phi_{n}}{\partial v_{j}}=m_{j}^{(n)}.$ (82)
@@
Appendix I Cyanobacteria experiments

### v3

@@ V Variational IB
$I(X_{t},H_{t})$ $\leq\int dx_{t}dh_{t}p(h_{t}|x_{t})p(x_{t})\log\frac{p(h_{t}|x_{t})}{\hat{p}(h_{t})}$ (S87)
$=\mathbb{E}_{X_{t}}D_{\text{KL}}(p(h_{t}|x_{t})\lVert\hat{p}(h_{t})).$ (S88)
If $p(h_{t}|x_{t})$ and $\hat{p}(h_{t})$ are chosen properly, the Kullback-Leibler can be expressed analytically which enables gradients to be effectively computed. As in Ref. [41], we take a Gaussian ansatz for $p(h_{t}|x_{t})$ and let the marginal $\hat{p}(h_{t})$ be a spherical unit-variance Gaussian. More concretely, encoded variables $H_{t}$ are sampled from $p(h_{t}|x_{t})$ by computing
$h_{t}=f_{W}(x_{t})+\sigma_{W}(x_{t})\eta,$ (S89)
@@
To bound the entire loss from above, we must bound $I(X_{t+\Delta t},H_{t})$ from below. We do this using the noise-contrastive estimate of the mutual information introduced in Ref. [111]. This recasts the problem as one of distinguishing samples from the distributions $p(x_{t+\Delta t}|h_{t})$ and $p(x_{t+\Delta t})$. Given a batch of $B$ pairs $(x_{t+\Delta t},h_{t})$ and one particular sampled value $h_{t}^{(i)}$, one asks what the probability is that a sampled value $x_{t+\Delta t}^{(j)}$ is from $p(x_{t+\Delta t}|h_{t}^{(i)})$ (is a positive sample) and not $p(x_{t+\Delta t})$ (is a negative sample). This probability is
$p(x_{t+\Delta t}^{(j)}=\text{pos}|h_{t}^{(i)})=\frac{\frac{p(x_{t+\Delta t}^{(j)}|h_{t}^{(i)})}{p(x_{t+\Delta t}^{(i)})}}{\sum_{k}^{B}\frac{p(x_{t+\Delta t}^{(k)}|h_{t}^{(i)})}{p(x_{t+\Delta t}^{(k)})}},$ (S90)
@@
$\mathbb{E}[-\log p(x_{t+\Delta t}^{(j)}=\text{pos}|h_{t}^{(i)})],$ (S91)
where the expectation is taken over indices $i$ and $j$, is closely related to the mutual information $I(X_{t+\Delta t},H_{t})$: in the limit of infinite samples $B\rightarrow\infty$, this quantity is given by
$-\int dx_{t+\Delta t}dh_{t}\,p(h_{t})p(x_{t+\Delta t}|h_{t})\log p(x_{t+\Delta t}=\text{pos}|h_{t})=\log B-I(X_{t+\Delta t},H_{t}).$ (S92)
If one had access to the probabilities $p(x_{t+\Delta t}^{(j)}=\text{pos}|h_{t}^{(i)})$ appearing in Eq. S91, the mutual information could thus be easily determined. One attempts to estimate these probabilities by introducing a variational ansatz $f(x_{t+\Delta t},h_{t})$ to approximate the density ratio $\frac{p(x_{t+\Delta t}|h_{t})}{p(x_{t+\Delta t})}$. Typically this $f$ is represented by a neural network. One can then obtain a bound on the mutual information by minimizing
$\mathcal{L}_{B}=\mathbb{E}\left[-\log\frac{f(x_{t+\Delta t}^{(i)}|h_{t}^{(i)})}{\sum_{k}^{B}f(x_{t+\Delta t}^{(k)}|h_{t}^{(i)})}\right],$ (S93)
@@
$I_{\text{NCE}}(X_{t+\Delta t},H_{t})=\log B-\mathcal{L}_{B}\leq I_{\text{true}}(X_{t+\Delta t},H_{t}).$ (S94)
@@
$\mathcal{L}_{\text{VIB}}=D_{KL}(p(H_{t}|X_{t})\lVert\hat{p}(H_{t}))-\beta I_{\text{NCE}}(X_{t+\Delta t};H_{t}),$ (S95)
@@
The VIB loss Eq. S95 is very similar to the loss function for a time-lagged $\beta$-variational autoencoder ($\beta$-VAE) [112], which have a form $\mathcal{L}=D_{KL}-\mathcal{L}_{\text{rec}}$, where the reconstruction term $\mathcal{L}_{\text{rec}}$ measures the deviation from the true future state $X_{t+\Delta t}$ and the reconstruction from the latent variable $H_{t}$. In the VIB, this term is replaced by the mutual information $I_{\text{NCE}}$: rather than searching for a latent variable which can reconstruct the While a standard $\beta$-VAE with mean squared error (MSE) loss finds a latent variable that can reconstruct the full state at a time $\Delta t$ in the future, the VIB merely tries to reconstruct the statistics of $X_{t+\Delta t}$ conditioned on $X_{t}$.
@@
$\dot{\theta}$ $=\omega.$ (S96)
Equation (S96) is the normal form (in polar coordinates) for dynamics near a Hopf bifurcation [118]. For $\mu>0$ this system exhibits a circular limit cycle of radius $\sqrt{\mu}$ (Fig. S8b). In Section B we show that this system has infinite purely imaginary eigenvalues $\lambda_{n}=in\omega$ and that as a result, encoders which exactly encode the angle coordinate, $p(h|r,\theta)\propto\delta(h-\theta)$, are solutions of the IB optimization problem.
The exact IB calculation breaks down for perfectly deterministic dynamics, hence we add a small amount of white noise to the dynamics. While this slightly perturbs the spectral content of the transfer operator, we still find our results to be consistent with the those expected for the deterministic case (see Section B). As we increase the size of the encoding alphabet $N_{H}$, the encoder partitions space by finer and finer angular wedges. The learned encoding is invariant under permutations of the encoded symbols. Upon reordering, we find that, for large alphabets $N_{H}\gg 1$, the encoder indeed approximates $p(h|r,\theta)\propto\delta(h-\theta)$ as expected (Fig. S8c).
Using VIB, we learn a continuous $h_{t}$ which can be computed directly from samples of the state variable. The encoding $h_{t}$ learned by VIB closely approximates the angle coordinate (Fig. S8e-f) and is nearly uncorrelated with $r$ (inset). This shows that our mathematical results illustrated for exact IB in Sec. IV hold also in the approximate framework of VIB.
@@
IB is an inherently probabilistic framework. To handle deterministic dynamics, we inject noise by introducing a stochastic sampling scheme. This step is necessary only for deterministically simulated dynamics; in our experience experimental data is sufficiently noisy on its own. Concretely, rather than taking as our IB relevance variable $Y=X_{t+\Delta t}$, we take $Y=X_{t+\Delta t+\eta}$ where $\eta$ is a random uniformly-distributed time shift. Despite using a different relevance variable, this yields essentially the same optimal encoding as (S47), where the eigenvalue $\text{e}^{\lambda_{n}\Delta t}$ is replaced by $\int d\eta\,p(\eta)\exp(\lambda_{n}(\Delta t+\eta))$. Crucially, the encoder retains its dependence on the transfer operator eigenfunctions $\phi_{n}(x)$ as before. In general, this need not be the case: selecting a new relevance variable changes the IB objective and will generically lead to a different encoder. Our choice of stochasticity which we introduce to the dynamics is chosen in such a way to preserve the form of the encoder.
Figure S9: VIB applied to the Lorenz-63 system (a) Illustration of the function found by VIB when allowed to learn a one-dimensional encoding. (b-c) Learned functions shown projected onto xyx-y, left is in the limit of low compression, while right is for high compression. (d) True eigenfunction, computed using the GAIO library [119]. (e) Time series of $h$ in the high- and low-compression limits.
@@
We illustrate this with a prototypical example of deterministic chaotic dynamics, the Lorenz system [106]. In the steady state, the state variable of the Lorenz system resides on a chaotic attractor consisting of two “lobes” which encircle unstable fixed points (Fig. S9a). Eigenvalues $\lambda_{n}$ as well as right eigenvectors $\phi_{n}(x)$ of the Perron-Frobenius operator were computed numerically using the Ulam method. The correspondence between $\phi_{1}(x)$ and $h$ disappears as $\beta$ is increased (Fig. S9b-d). The dynamics of $h$ also become notably less “slow”, and are instead nearly constant with occasional large jumps (Fig. S9e). The encoding variable $h$ learned by VIB is the slowest varying function of the state $\mathbf{x}$, decaying as $\exp(\lambda^{\text{Re}}_{1}t)$ where $\lambda^{\text{Re}}_{1}$ is the real part of the first subleading eigenvalue of the transfer operator (here Perron-Frobenius operator), which we show below in Fig. S11. This shows that in order for $h$ to be a valid slow variable, the compression term in Eq. (S95) is crucial. These results are consistent with those obtained for exact IB, where we showed that the encoder incorporates the slow modes at low beta (high compression).
VI VIB for fluid flow around a cylinder
Form of the eigenfunctions
$\phi_{n}[\mathbf{v}]=\langle\mathbf{v}(\mathbf{x}),\mathbf{m}^{(n)}(\mathbf{x})\rangle,$ (S97)
where $\mathbf{m}^{(n)}$ is the $n$-th mode (often referred to as a Koopman mode [27]) and angled brackets denote integration over space. The true eigenfunction and corresponding modes can be computed via dynamic mode decomposition (DMD) [15, 16], as described in the SI. The eigenfunctions for this system are in general complex, and come in conjugate pairs: $\phi_{2}(x)=\phi^{*}_{1}(x)$. In this situation any linear combination of $\phi_{1}$ and $\phi_{2}$ will decay at the same rate, and hence we expect to learn some arbitrary combination of the two dominant eigenfunctions, or equivalently a combination of the real and imaginary parts of $\phi_{1}$. We therefore take a two-dimensional encoding variable $h_{0},h_{1}]$, so that it can represent the full complex eigenfunction rather than only the real or imaginary part.
Our learned latent variables are oscillatory with the correct frequency as shown in Fig. 5b-c of the main text. A more stringent test is whether we are also learning the correct mode $\mathbf{m}^{(1)}$. From the true eigenfunctions, the modes can be extracted by computing the gradient
$\frac{\partial\phi_{n}}{\partial v_{j}}=m_{j}^{(n)}.$ (S98)
$\frac{\partial h}{\partial v_{j}}=m^{(\text{IB})}_{j}+g_{\text{res},j}(\mathbf{v}(\mathbf{x})),$ (S99)
where we have separated the part of the gradient which is independent of $\mathbf{v}$ from a residual part which is dependent on $\mathbf{v}$. If $h$ corresponds to the true eigenfunction, we expect that $\mathbf{m}^{(\text{IB})}$ is approximately equal to the Koopman mode $\mathbf{m}^{(1)}$, and that $\mathbf{g}_{\text{res}}$ is small. We indeed find this to be the case; Fig. 5d of the main text shows these gradients averaged over several instantiations of the neural network, which corresponds strongly to the true mode. Details concerning both the averaging procedure and the residuals $\mathbf{g}_{\text{res}}$ can be found in the SI. This shows that variational IB not only recovers the essential oscillatory nature of the dynamics, but does so by learning the correct slowly varying functions of the state variable given by the adjoint transfer operator eigenfunctions.
@@ VI VIB for fluid flow around a cylinder
Here we provide some additional details on the computation used to generate components of Fig. 5 of the main text. The true Koopman modes were computed using dynamical mode decomposition (DMD) [15, 16, 120], also described below in Section X. DMD attempts to find a finite dimensional approximation of the Koopman operator using $n$ snapshots of the system’s state $x\in\mathbb{R}^{d}$ which are assembled into a data matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$. One then attempts to find a linear evolution operator $K$ which propagates the state forward in time
@@
$\phi_{n}[\mathbf{v}]=\langle\mathbf{v}(\mathbf{x}),\mathbf{m}^{(n)}(\mathbf{x})\rangle,$ (S100)
where $\mathbf{m}^{(n)}$ is the $n$-th “Koopman mode” and angled brackets denote integration over space. To compute these modes and the corresponding transfer operator spectrum, we use dynamic mode decomposition (DMD; see SI Section VI) [15, 16].
We allow VIB to learn a two-dimensional encoding variable $h_{0},h_{1}]$, so that it can learn the complete first eigenfunction rather than only the real or imaginary part. For the purposes of comparing our learned variable with $\phi_{1}$, we construct a complex $h=h_{0}+ih_{1}$ out of the two learned components.
To understand which function has been learned by the neural network, we check whether the learned functions $h_{i}[\mathbf{v}]$ are of the form Eq. (S100). This can be done by examining gradients of the network with respect to the input field
$\frac{\partial h}{\partial v_{j}}=m^{(\text{IB})}_{j}+g_{\text{res},j}(\mathbf{v}(\mathbf{x}))$ (S101)
@@
$\frac{\partial\phi_{n}}{\partial v_{j}}=m_{j}^{(n)}.$ (S102)
@@
VII Chaotic systems
Chaotic systems exhibit many features which make them “difficult” to forecast. For example, they reside on strange attractors with fractal structure; even infinitesimal shifts in initial conditions lead to large deviations in a short time; and they are “mixing”, meaning that correlations decay to zero in time which is typically characteristic of stochastic systems. Although the goal of IB is compression (and not forecasting) we ask how the IB-optimal encoding reflects these properties. We do this in the contexts of the Lorenz-63 system [106], a hyperchaotic system [121], and the Kuramoto-Sivashinsky system [122, 123].
Fractal attractors
A unique feature of strange attractors is their fractal structure. While this is an interesting theoretical property, it is difficult to estimate from data in practice. The dominant methods for estimating fractal dimension from a time-series are little changed from those introduced nearly 50 years ago [124]. It is not known how to estimate the Hausdorff dimension of an attractor [125]; one typically relies on an estimate of the correlation dimension instead, which in many cases is close to the Hausdorff dimension [126].
One difficulty in the estimation of correlation dimension is that it requires the computation of distances between states of the system. In high-dimensional systems the effect of noise on measures of distance can become large: the ratio of distances between a point to its nearest and farthest neighbors approaches 1 [127, 128], which can lead to large corrections to the correlation dimension [129]. We will show that the encoding learned by VIB preserves the fractal dimension of the full system. This opens the door to using it in conjunction with denoising strategies to reliably determine the fractal dimension of very high-dimensional systems.
We consider two systems, a driven one-dimensional sine-Gordon (sG) equation and a one-dimensional Kuramoto-Sivashinsky (KS) equation. Both systems are described by time-evolving partial differential equations, and are hence high-dimensional. However, they exhibit low-dimensional dynamics. The sine-Gordon system is quasiperiodic and can thus be described by dynamics on a torus, while the KS system is chaotic and evolves on a low-dimensional strange attractor. We show in Fig. S12 and Fig. 6 of the main text that the fractal dimension of the respective embedded attractors is consistent with estimates of the full attractors. We find that the fractal dimension of attractor embedded in latent space coincides with that of the full system, which is much more computationally-expensive to estimate. This suggests that in the case where the true fractal dimension is unknown, it can be recovered by looking only at the dynamics in the encoding space.
Mixing dynamics
In terms of the transfer operator spectrum, a system is mixing if $P$ has only $\lambda=1$ in its point spectrum [130, 131], and otherwise consists only of a continuous spectrum (please see our comment below on “continuous spectra”). A more experimentally-relevant property of mixing systems is that correlations decay to zero $\langle A(t)B(0)\rangle-\langle A\rangle\langle B\rangle\overset{t\rightarrow\infty}{\rightarrow}0$. In this context, the primary quantity of interest is the subleading eigenvalue of the transfer operator spectrum, as this bounds the rate of decay [132]. Here we ask whether the latent variables learned by VIB, which approximate the subleading eigenfunction of the transfer operator, enable us to estimate mixing rates. As a demonstration, we consider the chaotic Lorenz-63 system as well as a hyperchaotic system governed by the equations of motion [121]:
$\dot{x}$ $=\alpha(y-x)+w$
$\dot{y}$ $=x(\beta-z)-y$
$\dot{z}$ $=xy-\gamma z$
$\dot{w}$ $=\delta w+\frac{1}{2}y^{2}-xz$
where $(\alpha,\beta,\gamma,\delta)=(10,28,2.7,2.2)$.
To ensure that our latent variable isolates the first eigenfunction and does not include further subleading eigenfunctions, we restrict $h$ to be one-dimensional and take a small compression parameter $\beta$. The results of the learned encoding are shown in Fig. S11a,d. The correlations of the latent variable decay on a slower time-scale than single components of the system Fig. S11b,e. For these relatively low-dimensional systems, we can independently approximate the dominant eigenvalue $\lambda_{1}$ using an Ulam approach, which we do using the software package GAIO [119]. We find good agreement between these eigenvalues and the decay rate estimated from a fit to the correlation function $C(\tau)=\langle h(t+\tau)h(t)\rangle-\langle h(t)\rangle^{2}$ where averages are taken over time Fig. S11c,f.
Figure S11: Mixing in chaotic systems
By finding the dominant eigenfunction of the transfer function, one can estimate the mixing rate of dynamical systems which we demonstrate on (a-c) the Lorenz system and (d-f) a 4-dimensional hyperchaotic system . (a) Attractor of the Lorenz system colored by the learned encoding value. (b) Time autocorrelations for the encoding $C_{hh}(\tau)=\langle h(t)h(t+\tau)\rangle_{t}-\langle h(t)\rangle_{t}^{2}$ (black) compared to the mixing rate computed by estimating the full transfer operator (blue; see text). The encoding variable decays over a much slower time scale than the state variable $x$ (gray). (c) Mixing rate estimated by fitting the blue curve in (b) to an exponential. We show the fitted real and imaginary parts of the decay rate separately. The true values are given by dashed lines of the respective color. (d-f) Show the analogous results for the hyperchaotic system mentioned in the text. Note that the full state is four-dimensional, so that in (d) we are only showing a projection onto three selected components.
Exponential divergence of nearby trajectories
Exponential divergence of nearby trajectories is a fundamental issue when attempting to forecast chaotic systems [133]. For that reason, modern methods typically avoid training neural networks to forecast dynamics using mean-square error alone, but rather include higher-order information about the dynamics, such as Jacobians [134] or place emphasis instead on preserving long-term statistics of the system [135, 98]. However, the goal of our work is not to predict the full dynamics: neither at the level of the trajectories, nor at the statistical level. Instead our goal is to find an appropriate dimensionally-reduced representation of the state of the system.
In Fig. S12a we consider the Kuramoto-Sivashinsky model, which governs the chaotic evolution of a one-dimensional continuous field [122, 123]. Though the system itself is high-dimensional (depending on the coarseness of discretization of the continuous field) it lives on a low-dimensional attractor (Fig. S12b) [136, 137]. We see that the top Lyapunov exponent, which characterizes the exponential divergence of nearby trajectories, is preserved by VIB when the dimensionality of the latent space is sufficient to fully encode the system’s attractor Fig. S12c. This fact is not necessarily surprising; as long as the encoding is a smooth transformation of points on the system’s attractor, this will generally be true. To see this, recall the definition of the top Lyapunov exponent [138]
$\lambda_{x}=\lim_{t\rightarrow\infty}\frac{1}{t}\log\lVert\delta x(t)\lVert$
where the small perturbation evolves according to $\dot{\delta x}_{i}=\frac{\partial f_{i}}{\partial x_{j}}\delta x_{j}$, with $f$ governing the system’s evolution $\dot{x}=f(x)$. Under a transformation $y=g(x)$, a small perturbation $\delta x$ yields the perturbation $\delta y_{i}=\frac{\partial g_{i}}{\partial x_{j}}\delta x_{j}$, or more concisely, $\delta y=D_{x}g\cdot\delta x$ Hence, the Lyapunov exponent estimated using the $y$ variables is given by
$\lambda_{y}$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\log\lVert\delta y(t)\lVert=\lim_{t\rightarrow\infty}\frac{1}{t}\log\lVert(D_{x}g)\delta x(t)\lVert$
We assume that the attractor in $x$ is bounded, and that the transformation $g$ is sufficiently smooth and invertible so that the attractor in $y$ is also bounded and $D_{x}g\leq C$ for some constant $C$. In addition, we will use the fact that $\lVert Ax\lVert\geq\sigma_{\text{min}}\lVert x\lVert$ where $\sigma_{\text{min}}$ is the minimal singular value of $A$. The Lyapunov exponent can consequently be bounded
$\lim_{t\rightarrow\infty}\frac{1}{t}\log\sigma_{\text{min}}\leq\lambda_{y}-\lambda_{x}\leq\lim_{t\rightarrow\infty}\frac{1}{t}\log C$
Assuming $\sigma_{\text{min}}$ and $C$ are both finite, in the limit this implies $\lambda_{y}=\lambda_{x}$.
This result breaks down when the transformation is non-invertible, in which case we have $\sigma_{\text{min}}\rightarrow 0$. So, while we expect the Lyapunov exponents to be captured as long as the encoding recovers the full attractor, for smaller dimensions the exponents may differ. This effect is demonstrated in Fig. S12d. The inset shows the evolution of a small perturbation $\langle\lvert\mathbf{\delta h}\rvert(t)\rangle$ over time, where the average is taken over $\mathcal{O}(10^{4})$ initial conditions. Different lines of the same color denote different instantiations of the same neural network (cross validation). For small dimensions, points which are distinct in the full system get mapped close together in the embedding space. In the case that they come from different trajectories, they may move apart very quickly (and could in principle even move in the opposite direction). This explains the rapid growth observed for small dimensions. For sufficiently high dimensional embeddings the exponent approaches the true value $\lambda\approx 0.048$, suggesting that the network has learned the full attractor. This demonstrates that the latent representation found by VIB retains the essential dynamical features of the full system.
Continuous spectra
In our work we take the perspective that noise is an unavoidable and essential feature of the system we aim to describe, which is reflected in our choice to invoke information theory and our focus on stochastic dynamical systems. While continuous spectra are relevant for the study of deterministic chaotic systems [27], in practice a system’s essential spectrum will shrink due to the presence of noise [50, 33]. This does not mean that continuous spectra are meaningless, just that the accompanying singularities are “smoothed.” Consider for example a Brownian particle in a double-well potential, as shown in the main text. In the deterministic limit, the system exhibits a continuous spectrum at the bifurcation $\mu=0$ which is associated with critical slowing down and the algebraic decay of correlations [49]. However, the presence of noise renders the spectrum discrete for all values of $\mu$. While the system still “slows down” near the deterministic bifurcation point, the massive qualitative change in the system’s dynamics is smoothed over [49].
Figure S12: Fractal dimension of VIB encodings
(a) Fractal dimension of the embedded Kuramoto-Sivashinsky attractor with varying encoding dimension. We estimate fractal dimension here by computing the correlation dimension [125, 139]. The gray dashed line shows the upper bound on the fractal dimension given by the Kaplan-Yorke conjecture [140, 141]. Green dashed line denotes the dimension of the attractor approximated for the full system, which lives in an $\mathbb{R}^{N}$, $N=128$ discretization of the system’s domain. b: Fractal dimension of the embedded system as a function of encoding dimension. Dashed line shows the fractal dimension of the system measured using the full state. c: Mutual information between the embedding and the future state, as a function of encoding dimension. d: At sufficiently high embedding dimension, encodings preserve the system’s Lyapunov exponent. Dashed line shows the true exponent of $\lambda\approx 0.048$. Inset shows the distance $\lvert\delta h\rvert$ as a function of time for points which start close together (initial distance is normalized to one for comparison).
VIII Cyanobacteria experiments